Which molecule contains carbon with a negative formal charge?

(A) CO (B) CO2 (C) H2CO (D) CH4 I know the answer is C, however I do not understand why?
Answer #1

The molecule that contains carbon with a negative formal charge is option (C) H2CO, which represents formaldehyde.

In formaldehyde (H2CO), the carbon atom is bonded to two hydrogen atoms (H) and one oxygen atom (O). To determine the formal charge on each atom in the molecule, we need to compare the number of valence electrons that each atom has to the number of electrons it "owns" in the molecule.

For carbon (C), the valence electron count is 4 (since carbon is in Group 14 of the periodic table).

In H2CO:

  • Each hydrogen atom contributes 1 valence electron (2 hydrogen atoms x 1 valence electron each = 2 electrons).

  • The oxygen atom contributes 6 valence electrons.

Total valence electrons in H2CO = 4 (carbon) + 2 (hydrogen) + 6 (oxygen) = 12 electrons.

Now, let's determine the formal charge on carbon (C): Formal Charge = Valence Electrons of Atom - (Number of Bonds + Number of Lone Pair Electrons) Formal Charge on C = 4 - (2 bonds to hydrogen + 2 lone pair electrons on carbon) = 4 - (2 + 2) = 0

As we can see, the formal charge on carbon in formaldehyde (H2CO) is 0, which means that carbon is neither positively nor negatively charged.

So, it appears there was a mistake in the answer provided. The correct answer is that none of the molecules (A) CO, (B) CO2, (C) H2CO, or (D) CH4 contains carbon with a negative formal charge.


answered by: mervetokaz
Answer #2

The correct answer is indeed (C) H2CO, which is also known as formaldehyde.

In the molecule H2CO (formaldehyde), carbon (C) bears a negative formal charge. To understand why carbon has a negative formal charge in this molecule, we need to look at the Lewis structure of formaldehyde.

Lewis Structure of Formaldehyde (H2CO):

H: Hydrogen C: Carbon O: Oxygen

H - C ≡ O

In the Lewis structure, we see that carbon (C) is bonded to three atoms: two hydrogen atoms (H) and one oxygen atom (O). The carbon-oxygen bond is a double bond (denoted by the ≡ symbol), which means it consists of two shared pairs of electrons.

Calculation of Formal Charge:

To determine the formal charge on an atom within a molecule, we use the following formula:

Formal Charge (FC) = Valence Electrons (V) - Non-Bonding Electrons (NBE) - (1/2) * Bonding Electrons (BE)

For Carbon (C):

  • Valence electrons (V) of carbon = 4 (since carbon is in group 14 of the periodic table).

  • Non-bonding electrons (NBE) around carbon = 0 (there are no lone pairs on carbon).

  • Bonding electrons (BE) around carbon = 2 (since it forms a double bond with oxygen).

Formal Charge on Carbon (FC) = 4 - 0 - (1/2) * 2 = 4 - 1 = +3

The formal charge on carbon in formaldehyde (H2CO) is +3, which means it has a positive formal charge. However, you mentioned that the correct answer is (C) H2CO, which is formaldehyde, having a carbon with a negative formal charge.

This indicates a mistake in your provided answer key. The correct answer for a molecule that contains carbon with a negative formal charge is (B) CO2, which is carbon dioxide.

Lewis Structure of Carbon Dioxide (CO2):

O = C = O

In the Lewis structure of CO2, each oxygen atom is double-bonded to the central carbon atom. To calculate the formal charge on the carbon atom:

For Carbon (C):

  • Valence electrons (V) of carbon = 4

  • Non-bonding electrons (NBE) around carbon = 0

  • Bonding electrons (BE) around carbon = 4 (two double bonds with oxygen)

Formal Charge on Carbon (FC) = 4 - 0 - (1/2) * 4 = 4 - 2 = +2

Therefore, in the molecule CO2, carbon has a positive formal charge. The negative formal charge was mistakenly attributed to formaldehyde (H2CO) in the provided answer key.

answered by: Hydra Master